you are lying 120 ft away from a tree that is 50 feet tall. you look up at the top of the tree. approximatley how far is your head from the top of the tree in a straight line
Answer by checkley79(3341)

A^2+B^2=C^2
120^2+50^2=C^2
14,400+2,500=C^2
16,900=C^2
C=SQRT16,900
C=130 FEET IS THE DISTANCE TO THE TOP OF THE TREE.